# itemizestditemize

### v1

@@ II Overview of Generative Predictive Control
GPC-RANK and GPC-OPT can be combined. We can sample $K$ action proposals from ${\cal P}(\cdot)$ and for each proposal perform reward maximization according to ((ii)), leading to $K$ optimized action chunks. We pick the optimized action chunk with highest reward. GPC-RANK+OPT can be viewed as solving the optimization (3) from multiple initializations [50], and requires $K\times M$ evaluations of the world model.
We provide a succinct summary of GPC in Algorithm 1.
@@
5// Predictive world modelingIII)
@@ 18      Set $a^{(k)}_{t:t+T}=\hat{a}^{(M)}_{t:t+T}$;
Remark 1 (Connection with Offline RL).
It is worth noting that the composition of the reward model ${\cal R}(\cdot)$ and the world model ${\cal W}(\cdot)$ can be identified as the so-called Q-value function in offline RL, as they predict the reward of an action chunk $a_{t:t+T}$ given the current information $I_{t}$. However, there are three subtle differences between GPC and offline RL. (a) GPC learns an explicit dynamics model instead of just approximating the Q-value function (this is similar to model-based offline RL). (b) The composition of ${\cal R}(\cdot)$ and ${\cal W}(\cdot)$ provides a Q-value function for “action chunks” while the usual Q-value function involves a single action. (c) GPC learns the dynamics model using diverse trajectories (without rewards), instead of only expert demonstrations. In §IV, we compare GPC with an offline RL method adapted from single action to action chunk and show that GPC outperforms offline RL.
The above overview of GPC highlights the importance of the world model in both ranking and optimizing action proposals. Indeed, the concept of a learned world model dates back to early pioneering works on deep visual foresight [20]. In the next section, we will show that modern diffusion models are the key enabler for learning (near) physics-accurate visual world models that were not possible before.
Figure 2: Illustration of planar pushing.
III World Model Learning
Recall that the world model takes as input $(I_{t},a_{t:t+T})$ and outputs $I_{t+1:t+T+1}$. We will briefly present the world model design when $I_{t}=x_{t}$ contains low-dimensional states (§III-A) and focus on a diffusion-based world model when $I_{t}=o_{t-H:t}$ contains high-dimensional images (§III-B).
@@ III World Model Learning
Planar pushing. Before getting started, since we will focus on planar pushing in our experiments §IV, we briefly introduce the problem setup. As shown in Fig. 2, the planar pushing task involves pushing an object (e.g., a T block) to align with the target using a pusher. We assume the target position is fixed and known, and hence we do not need to predict it. The state of the object is fully described by its 2D coordinates $p=(p_{x},p_{y})$ and 1D angle $\theta$. The action space is $a\in\mathbb{R}^{2}$ containing pushing velocities in the $x$ and $y$ directions. We choose the planar pushing task for two reasons. First, it is easy to understand and simple to visualize. Second, it is also a long-standing difficult task because it is contact-rich. Depending on whether the pusher makes or breaks contact and whether the contact is sticking or sliding, the dynamics of the task involves multiple different modes and requires intricate physics reasoning [26]. This makes planar pushing a both elementary and challenging task for world model learning.
III-A State-based World Modeling
@@
For state-based world modeling, with $x=(p,\theta)$ being the state of the object, we learn a single-step dynamics model:
$x_{t+1}=f_{\text{state}}(x_{t},a_{t}).$ (6)
@@
for both the input and output of $f_{\text{state}}$ in (6) (we normalize the output of $f_{\text{state}}$ to satisfy $\sin^{2}(\cdot)+\cos^{2}(\cdot)=1$). We also found it beneficial to have two separate networks, one predicting the position and the other predicting the rotation. In our implementation, for each of the network, we used a six-layer MLP with hidden dimension $128$.
During testing, we need to predict $T$ steps into the future given a sequence of actions $a_{t:t+T}$. We do so by recursively applying the single-step predictor (6). As we will show in §IV, this simple MLP-based single-step predictor is sufficient for robust online planning.
III-B Vision-based World Modeling
State-based world modeling requires accurate state estimation, which is possible in a lab environment with infrastructure such as AprilTags [46] or motion capture but can be challenging in the open environment. Therefore, it is desired to directly learn visual dynamics, i.e., a visual world model.
Diffusion-based visual world modeling. Motivated by the success of using diffusion models for image generation, we design a visual world model based on conditional diffusion. Recall that the input to the world model is $I_{t}=o_{t-H:t}$ (the sequence of past images) and $a_{t:t+T}$, and the output is $I_{t+1:t+T+1}:=o_{t+1:t+T+1}$ (a sequence of future images). Similar as the state-based world modeling, we design the visual world model as a recursive application of single-step image predictors. We denote the single-step predictor as
$o_{t+1}=f_{\text{vision}}(o_{t-H:t},a_{t}),$ (8)
where ${\cal T}$ is the total number of denoising steps and the output of $f_{\text{vision}}$ is $o_{t+1}=o_{t+1}^{0}$. In (9), $\tau$ is the denoising step index and $D_{\phi}$—the “denoiser”—is a neural network with weights $\phi$. We use the same architecture for $D_{\phi}$ as in [5], which contains convolutions, action embedding, and a U-Net, as illustrated in Fig. 3. The denoiser $D_{\phi}$ is trained by adding random noises to the clean images and then predicting the noise.
Two-phase training. To improve the accuracy and consistency of visual world modeling, we train it in two phases. In phase one, we train only the single-step image predictor (8), i.e., with supervision only on a single image $o_{t+1}$. In phase two, we recursively apply the single-step predictor (8) for $T$ times to obtain a sequence of future images $o_{t+1:t+T+1}$, and jointly supervise them with groundtruth images.
We use observation horizon $H=4$ in the visual world modeling, and ${\cal T}=3$ diffusion steps in $f_{\text{vision}}$.
Remark 2 (Freeze the Noise).
It should be emphasized that the single-step image predictor $f_{\text{vision}}$ in (8) is “stochastic” where the stochasticity comes from the randomness in the initial noise $o_{t+1}^{{\cal T}}$. In classical generative modeling, stochasticity is desired because we want the model to generate diverse images. However in our paper, we want to use the diffusion-based world model for control. That is, we want to understand the impact of the “action” on the future, rather than the impact of the random “noise” on the future. Therefore, we freeze the noise $o_{t+1}^{{\cal T}}$ to be zero at inference time, which makes the world model deterministic—outputing the most likely future predictions (because $o_{t+1}^{{\cal T}}=0$ attains the peak probability in the Gaussian distribution). In fact, without freezing the noise, GPC-OPT does not work because the gradient of the reward with respect to the actions (cf. ((ii))) would be “contaminated” by the noise and thus does not guide online planning.
IV Experiments
In this section, we focus on evaluation of the GPC framework on planar pushing and comparison to baseline behavior cloning. As mentioned in §III, we choose planar pushing because it is an elementary but also quite challenging task due to the complicated contact and collision dynamics involved.
IV-A State-based Planar Pushing in Simulation
Policy. We first train a state-based diffusion policy ${\cal P}(\cdot)$ [15] from which we can draw action proposals, as explained in §II. We use $500$ expert demonstrations to train the policy network for $300$ epochs with AdamW optimizer of learning rate $10^{-4}$ and weight decay $10^{-6}$. We use observation horizon $H=2$, prediction horizon $T=16$, and action horizon $8$. [^2] This means the policy is executed in a receding horizon fashion. The policy predicts $16$ steps of future actions and only $8$ of them are executed in the environment, which creates implicit feedback.
@@ IV-A State-based Planar Pushing in Simulation
World model. Then we need the world model to predict the future consequences of each action proposal. As in §III-A, we use MLP networks to build a single-step state predictor and recursively apply it $T$ times to form the world model ${\cal W}(\cdot)$. As introduced in §II, to ensure the dynamics of the world model is learned from a richer distribution than the expert demonstrations, the training dataset of the world model is expanded with random exploration demonstrations which attempt to record more pusher-object interactions at different contact points. To be specific, we let human operators (who are different from the operator of the expert demonstrations) collect $7$ long random play demonstrations and 1 special long demonstration that randomly collides the object with the boundary wall. In addition, we also generate $7$ sets of “noisy expert demonstrations” where the original expert actions are contaminated with different scales of Gaussian noise.
@@
Results. In Table I, we study how different variations of $K$ and $M$ affect the performance. When $K=1$ and $M=0$, this is the baseline state-based diffusion policy trained from behavior cloning [15]. In GPC-RANK, if we increase the number of action proposals to $K=100$, the score increases to around 0.9, which is about $10\%$ increase from the baseline. Notably, due to parallelization inside GPC-RANK, the runtime does not increase when $K=100$. We also tried larger $K$, such as $1000$ and $5000$, and the performance further increases to $0.934$. Note that in the last column of Table I, we provided a “performance upper bound” which is obtained by using the groundtruth simulator (instead of the learned world model) to rank action proposals at $K=5000$. For GPC-OPT, if we sample one action proposal and perform gradient ascent from it (e.g., doing $30$ gradient steps), the score can be increased to $0.886$, which is also quite significant compared to the baseline. Then in GPC-RANK+OPT, we tried different combinations of $K$ and $M$. The improvement remains substantial, indicating the robustness of GPC and shows both sampling multiple proposals and gradient-based optimization consistently improve the performance.
In the first two columns of Table I, we present the results for another baseline of using Offline RL to rank and select actions. The Offline RL method follows that of [44] and is described in Appendix -A. Crucially, the Offline RL method only uses the expert demonstrations and does not leverage the random exploration data (because the random exploration data do not have labeled rewards). We can see that the performance of Offline RL is similar or even worse than the case of behavior cloning. A plausible reason for this is that the Offline RL method does not explicitly seek to learn the dynamics from the random explorations and hence the learned value function may be misleading for ranking action proposals.
Finally, in Table II, we present the results of using a world model trained without random exploration data. We observe that, when the world model is only trained from expert demonstrations, the performance of both GPC-RANK and GPC-OPT gets worse. This demonstrates the importance of using random exploration data to learn an accurate world model.
TABLE I: State-based Planar Pushing (Push-T) in Simulation. Score is measured by the IoU metric averaged over $100$ evaluation seeds. $K$ is the number of action proposals sampled from ${\cal P}(\cdot)$ while $M$ is the number of gradient steps performed to maximize reward.
@@
TABLE II: The importance of random exploration in world model learning (state-based Push-T).
IV-B Vision-based Planar Pushing in Simulation
Figure 4: Comparison of different visual world modeling techniques on two examples.
We then proceed to vision-based planar pushing with the same T-shape object.

### v2

@@ 2 Overview of Generative Predictive Control
GPC-RANK and GPC-OPT can be combined by sampling $K$ action proposals from ${\cal P}(\cdot)$ and applying reward maximization (4) to each, yielding $K$ optimized action chunks. We select the one with the highest reward. GPC-RANK+OPT effectively solves (3) from multiple initializations [26], requiring $K\times M$ world model evaluations. A summary of GPC is provided in Appendix A, with an analysis of its connection to offline RL in Appendix E.
The world model is crucial for ranking and optimizing action proposals. While early works such as deep visual foresight [27] explored learned world models, we next show that modern diffusion models now enable (near) physics-accurate visual predictions previously unattainable.
3 World Model Learning
The world model takes as input $(I_{t},a_{t:t+T})$ and predicts $I_{t+1:t+T+1}$. When $I_{t}$ represents a low-dimensional state, an MLP can be used to learn the dynamics, as demonstrated in Appendix C for the case of planar pushing. However, state-based world modeling relies on accurate state estimation, which is feasible in controlled lab environments with infrastructure such as AprilTags [28] or motion capture systems, but can be challenging in open or unstructured environments. Therefore, it is desirable to learn visual dynamics directly, i.e., to build a visual world model.
Diffusion-based visual world modeling. Motivated by the success of using diffusion models for image generation, we design a visual world model based on conditional diffusion. Recall that the input to the world model is $I_{t}=o_{t-H:t}$ (the sequence of past images) and $a_{t:t+T}$, and the output is $I_{t+1:t+T+1}:=o_{t+1:t+T+1}$ (a sequence of future images). We design the visual world model as a recursive application of single-step image predictors, which reads
$o_{t+1}=f_{\text{vision}}(o_{t-H:t},a_{t}),$ (6)
where ${\cal T}$ is the total number of denoising steps and the output of $f_{\text{vision}}$ is $o_{t+1}=o_{t+1}^{0}$. In (7), $\tau$ is the denoising step index and $D_{\phi}$—the “denoiser”—is a neural network with weights $\phi$. We use the same architecture for $D_{\phi}$ as [29], containing convolutions, action embedding, and a U-Net (Fig. 2). $D_{\phi}$ is trained by adding random noises to the clean images and then predicting the noise.
Two-phase training. To improve the accuracy and consistency of visual world modeling, we train it in two phases. In phase one, we train only the single-step image predictor (6), i.e., with supervision only from a single image $o_{t+1}$. In phase two, we recursively apply the single-step predictor (6) for $T$ times to obtain a sequence of future images $o_{t+1:t+T+1}$, and jointly supervise them with groundtruth images. We use observation horizon $H=4$ in the visual world modeling, and ${\cal T}=3$ diffusion steps.
Remark 1 (Freeze the Noise).
It should be emphasized that the single-step image predictor $f_{\text{vision}}$ in (6) is “stochastic” where the stochasticity comes from the randomness in the initial noise $o_{t+1}^{{\cal T}}$. In classical generative modeling, stochasticity is desired because we want the model to generate diverse images. However in our paper, we want to use the diffusion-based world model for control. That is, we want to understand the impact of the “action” on the future, rather than the impact of the random “noise” on the future. Therefore, we freeze the noise $o_{t+1}^{{\cal T}}$ to be zero at inference time, which makes the world model deterministic—outputting the most likely future predictions (because $o_{t+1}^{{\cal T}}=0$ attains the peak probability in the Gaussian distribution). In fact, without freezing the noise, GPC-OPT does not work because the gradient of the reward with respect to the actions (cf. (4)) would be “contaminated” by the noise and thus does not guide online planning.
4 Experiments
We evaluate GPC on (1) a state-based planar pushing task, (2) four vision-based simulation tasks, and (3) two real-world tasks. We found GPC consistently outperforms the behavior cloning baseline. Detailed descriptions of the tasks can be found in Appendix B.
4.1 State-based Planar Pushing in Simulation
We study the planar pushing task with the goal of pushing an object from an initial pose to a specified target pose, where the groundtruth pose of the object is available through a simulator. We firstly train a state-based diffusion policy ${\cal P}(\cdot)$ [7] from which we can sample multiple action proposals, as explained in §2. Then with both expert demonstrations and random exploration data, we utilize MLP networks to construct the state-based world model predicting the outcome states and define the reward with a registration loss between the target and predicted object states. Please refer to more training details in Appendix C.
In Table 1, we present results for GPC, the pure behavior cloning baseline ($K=1$ and $M=0$), and Offline RL baseline (more details in Appendix E). Clearly, all GPC-RANK, GPC-OPT, and GPC-RANK+OPT outperform the baselines. Notably, GPC-RANK+OPT approaches the performance of planning with a groundtruth simulator.
Table 1: State-based Planar Pushing (Push-T) in Simulation. Score is measured by the IoU metric averaged over $100$ evaluation seeds. $K$ is the number of action proposals sampled from ${\cal P}(\cdot)$ while $M$ is the number of gradient steps performed to maximize reward.
4.2 Vision-based Tasks in Simulation
Figure 3: World model prediction in GPC for simulation tasks.
Push-T Triangle Drawing Block Stacking Cube & Sphere Swapping Reward Learned Learned VLM Learned VLM VLM Behavior Cloning $0.642$ $0.724$ $0.724$ $0.912$ $0.912$ 0.68 GPC-RANK 0.739 0.775 0.755 0.988 0.987 0.73
Table 2: Scores for four vision-based tasks in simulation.
Figure 4: Comparison of different visual world modeling on two examples. (Highlights for comparison are boxed in blue).
We test on four vision-based simulation tasks named as Push-T, Triangle Drawing, Block Stacking, and Cube & Sphere Swapping. We train a vision-based diffusion policy ${\cal P}(\cdot)$ [7] using ResNet18 and UNet. For the world model, as in §3, we use a diffusion model to build a single-step image predictor and recursively apply it $T$ times to form the world model ${\cal W}(\cdot)$. Same as in §4.1, the training data for world model includes random explorations. Please see more training details in Appendix D.
Reward. Different from the state-based planar pushing, where future states are structured, the image-based predictions make it difficult to directly compute a reward signal. Therefore, we propose two ways to implement the reward function. (1) When a numerical reward can be defined, such as the registration loss between the T-block and the target frame in Push-T, or the distance between two cubes in Block Stacking, we train a reward predictor with a ResNet18 backbone followed by an MLP, mapping images to scalar rewards. The learned reward model $\mathcal{R}(\cdot)$ is differentiable and can be used for online planning. (2) For tasks where numerical rewards are tedious or infeasible to define, we propose leveraging VLM (ChatGPT-4o [30]) to select among action proposals in a zero-shot manner. After predicting outcomes with the world model, we prompt the VLM with the predicted images and the task goal, allowing it to choose the best action proposal.
Quality of visual world modeling. Fig. 3 illustrates the predicted images generated by our world model, demonstrating both visually compelling results and accurate predictions of object interactions and movements. Then we study whether our visual world modeling technique is better than some baselines. Particularly, we chose two baselines, one from the pioneering deep visual foresight [27], and the other from the more recent AVDC paper [15]. The world model in deep visual foresight leverages CNN and LSTM for future prediction, and we follow the Github implementation. [^1] https://github.com/Xiaohui9607/physical_interaction_video_prediction_pytorch AVDC is originally designed for video diffusion conditioned on language, and we modify its official implementation [^2] https://github.com/flow-diffusion/AVDC to condition on robot actions. Notably, AVDC is a multiple-step world model as it simultaneously diffuses future images together, unlike our recursive structure. Fig. 4 compares the quality of future predictions of three methods (ours, deep visual foresight, and AVDC) against the groundtruth in the Push-T task. We see that diffusion-based visual world modeling (ours and AVDC) clearly dominates the classical deep visual foresight. Between ours and AVDC, we can observe that the future predictions from our method are better aligned with the groundtruth.
Planning performance. We present the results of GPC-RANK and pure behavior cloning baseline in Table 2. With either a learned reward predictor ${\cal R}(\cdot)$ to pick from $100$ action proposals or VLM with $10$ candidates, GPC-RANK can significantly improve the pure behavior cloning baseline for all four vision-based tasks. Fig. 5 shows evaluation trajectories comparing GPC-RANK and baseline.
Figure 5: Evaluation Trajectories for Triangle Drawing and Cube & Sphere Swap.
Baseline Methods Method Offline RL ($K=100,M=0$) Behavior Cloning ($K=1,M=0$) CEM ($K=100,M=0$) CEM ($K=200,M=0$) Gradient Ascent ($K=1,M=25$) Score $0.620$ $0.642$ $0.174$ $0.202$ $0.131$
GPC Method GPC-RANK ($K=50,M=0$) GPC-RANK ($K=100,M=0$) GPC-OPT ($K=1,M=25$) GPC-RANK+OPT ($K=10,M=25$) Score 0.698 0.739 0.791 0.882
Table 3: Ablation Study for Vision-based Planar Pushing (Push-T) in Simulation. Score is measured by the IoU metric averaged over 100 evaluation seeds. $K$ is the number of action proposals and $M$ is the number of gradient steps.
Detailed comparison in planar pushing. We study the impact of $K$ and $M$ on the performance of GPC, its comparison with additional baselines, and the importance of random exploration.
Figure 6: Importance of random exploration in world model learning (vision-based push-T)
In Table 3, we study how different variations of $K$ and $M$ affect the performance of GPC. Since image operations require significantly more memory and computation than state-based inputs, we limit $K$ to the scale of hundreds in vision-based experiments. The results show that: (a) GPC-RANK improves performance by approximately $10\%$ over the diffusion policy baseline ($K=1$, $M=0$); (b) GPC-OPT yields a $15\%$ improvement; and (c) GPC-RANK+OPT achieves up to a $25\%$ gain. Table 3 also includes comparisons with four baselines: Offline RL [31], pure behavior cloning [7], the cross entropy method (CEM) [32], and pure gradient ascent [32]. These results indicate that neither behavior cloning, offline RL, nor standalone planning methods are sufficient to achieve high performance in complex vision-based tasks. In contrast, GPC, which unifies generative and predictive control, consistently outperforms these baselines.
In Fig. 6, we compare the performance of GPC-RANK and GPC-OPT for Push-T when trained with or without random exploration. The performance is increased by $10\%$ approximately when introducing random exploration data, which demonstrates its importance in achieving an accurate world model. Similar results for state-based task are presented in Appendix C.
4.3 Real-world Vision-based Tasks
We then test GPC on real-world tasks, involving push-T (and its variations with object collisions) and clothes folding. For the push-T tasks, rewards are computed using a learned reward predictor composed of a ResNet and MLP, trained with registration losses derived from expert demonstrations’ state information using AprilTags [28]. However for cloth folding, where states are harder to define, we design a progressing reward: starting at 00 for spread clothes and increasing with each successful grasp-and-move operation. Normalized progressing rewards for expert demonstrations are used to train the reward predictor for clothes folding.
For plain push-T, while the baseline model succeeds 5 out of 10 times, both GPC-RANK and GPC-OPT succeeds 7 out of 10 times. In a more challenging version of push-T with obstacles (letters "A" and "R") placed on the table, out of 5 test trials, the baseline succeeds 2 times, GPC-RANK succeeds 3 times, and GPC-OPT succeeds 4 times. For the clothes folding task, while the baseline model succeeds 3 out of 10 times, GPC-RANK succeeds 7 out of 10 times. Fig. 7 and Fig. 8 show the trajectories for these tests.
Despite the challenging nature of these real-world experiments—where dynamics are more complex due to collisions and tasks like cloth folding involve non-rigid bodies and require multiple steps—GPC still operates effectively. More trajectories and the quality of world model prediction for real-world tests are shown in Appendix F and https://computationalrobotics.seas.harvard.edu/GPC/.
Figure 7: Real-world tests for push-T with collision. In every test, top row shows trajectories of baseline model ($K=1,M=0$), middle row shows trajectories of GPC-RANK ($K=10,M=0$), and last row shows trajectories of GPC-OPT ($K=0,M=25$).
Figure 8: Real-world tests for clothes folding. Top row shows trajectories of baseline model ($K=1,M=0$), second row shows trajectories of GPC-RANK ($K=10,M=0$).
5 Conclusion
We presented generative predictive control (GPC), a learning-based control paradigm that integrates a generative policy with a predictive world model for online planning. While GPC shares strong conceptual ties with classical predictive control, namely identifying a dynamics model followed by online optimization, it introduces several novel insights. The new insights offered by GPC are that (a) modern conditional diffusion models provides a powerful tool to learn controllable action-conditioned visual world models; (b)VLM based method can be used to pick from imagined outcomes of visual world models, as an alternative to learned reward predictors; and (c) a generative policy trained from behavior cloning provides good warmstarts for online optimization that helps mitigate the challenges posed by nonconvex reward landscapes in online optimization. Through extensive experiments in both simulated and real-world settings, we demonstrated the robustness and versatility of GPC in both state-based and vision-based robotic manipulation.
6 Limitation
Two issues are worth highlighting. The first issue is about hallucination. One of the important issues of generative modeling is hallucination, and our visual world model is no exception. In particular, there are several cases where the world model predicts images of the future that violate rigid-body physics. For example, the objects T, A, and R may penetrate each other, as we show in Fig. 9. Interestingly, even though the predictions can violate physics, using such predictions for planning still leads to GPC’s superior performance. We believe learning a visual world model that strictly satisfies the laws of physics remains a challenging task, and strategies such as data scaling and baking in physics priors may help solve this challenge. For example, it may be possible to leverage the so-called world foundation models [33]—trained on Internet-scale videos—and finetune them for specific robotic tasks.
The second issue is associated with the computational complexity of using an expensive diffusion-based world model in online planning. It is worth noting that, although GPC dominates the baseline in terms of success rates, it is generally slower than the baseline due to predicting the future using the world model. A promising solution is to leverage diffusion distillation diffusion [34, 35], and faster diffusion solvers [36] for faster world model prediction.
Figure 9: Limitations of world model prediction.
Appendix A GPC Pseudocode
@@
5// Predictive world modeling3)
@@ 18      Set $a^{(k)}_{t:t+T}=\hat{a}^{(M)}_{t:t+T}$;
Appendix B GPC Task Descriptions
In this section, we describe the 4 simulation tasks and 2 real-world tasks including the variations.
B.1 Simulation tasks
@@
  • Push-T. This is a planar pushing task involving pushing an object (e.g., a T block) to align with the target using a pusher. We use the same simulator as in [7] with a T-shape object. We assume the target position is fixed and known, and hence we do not need to predict it. The low-dimensional state of the object is fully described by its 2D coordinates $p=(p_{x},p_{y})$ and 1D angle $\theta$. The action space is $a\in\mathbb{R}^{2}$ containing pushing velocities in the $x$ and $y$ directions. Push-T is a long-standing difficult task because it is contact-rich. Depending on whether the pusher makes or breaks contact and whether the contact is sticking or sliding, the dynamics of the task involves multiple different modes and requires intricate physics reasoning [37]. The performance score is evaluated by the IoU value between block and target.
  • Triangle Drawing. This is a simulation task built with ManiSkill simulator [38]. The robotic hand is holding a pen to trace the triangle pattern on the paper. This is a task having high requirement about vision since it needs to trace triangle appearing in different location and orientation. The performance score is evaluated by the collapsed pixels between drawing and pattern.
  • Block Stacking. This [39] is a classical manipulation task, aiming to stack one block onto another target block. This involves moving the block exactly above the target and and putting the block down. It is also built with the ManiSkill simulator [38]. The performance score is evaluated by the distance between two blocks.
  • Cube & Sphere Swap. This is also a simulation task built with ManiSkill simulator [38]. Initially, there is a cube in the box and a sphere outside of the box. It requires the robot arm to move the cube out of the box and then move the sphere into the box. This task is challenging because it requires a logical operation sequence. Moreover, it is a multi-object tasks, which is challenging for the world-model prediction. The performance score is evaluated by completion status: $1.0$ is fully finished, $0.5$ is the half completion (only cube is moved out), $0.0$ is not done.
B.2 Real-world tasks
  • Plain Push-T. Since planar pushing is a challenging and elementary task, we also include the push-T task and its variations in real-world experiments. The pusher is attached to the end-effector of a Franka Panda robotic arm to push a T block into the target position. An overhead Intel RealSense camera is used to provide image observations to the pusher. To train the reward model, we need to obtain block information as training data so we attached three AprilTags [28] to the T-block, which enables accurate pose estimation from image frames even in the case of occlusion. To collect real-world expert demonstrations, we follow Chi et al. [7] and use a SpaceMouse to teleoperate the robotic arm to push the T block. A and R blocks are the possible objects that T may collide into during pushing.
  • Push-T with obstacle A. To make task more challenging, we put an object of shape “A” on top of the target position so that when pushing T, it will inevitably collide with object A. The world model here not only needs to learn the dynamics for pushing T block, but also capture the dynamics about interactions between T-block and A-block.
  • Push-T with obstacle A and R. Similarly, we put an object of shape “A” and “R” on top of the target position so that when pushing T, it will inevitably collide with object A and R. Here the dynamics is even more complex.
  • Push-T in unseen scenarios. Here we use the policy and world model trained with push-T, push-T colliding A, push-T colliding A & R expert demos. But we perform the test to push T when an object of letter “R” is on top of the target. This situation has never been seen in any of the training data.
  • Clothes Folding. We put a piece of clothes in random position and orientation on the table and ask the Franka hand to fold the clothes. It involves three steps: fold the left arm, fold the right arm, and fold the cap. Moreover, this is a task about operating on non-rigid body, making it quite challenging for both policy and world model.
Appendix C State-based Planar Pushing
Policy. We first train a state-based diffusion policy ${\cal P}(\cdot)$ [7] from which we can draw action proposals, as explained in §2. We use $500$ expert demonstrations to train the policy network for $300$ epochs with AdamW optimizer of learning rate $10^{-4}$ and weight decay $10^{-6}$. We use observation horizon $H=2$, prediction horizon $T=16$, and action horizon $8$. [^3] This means the policy is executed in a receding horizon fashion. The policy predicts $16$ steps of future actions and only $8$ of them are executed in the environment, which creates implicit feedback.
@@ Appendix C State-based Planar Pushing
World model. Then we need the world model to predict the future consequences of each action proposal. For state-based world modeling, since we focus on planar pushing tasks here, we define with $x=(p,\theta)$ being the state of the object, then we learn a single-step dynamics model:
$x_{t+1}=f_{\text{state}}(x_{t},a_{t}).$ (A1)
@@
for both the input and output of $f_{\text{state}}$ in (A1) (we normalize the output of $f_{\text{state}}$ to satisfy $\sin^{2}(\cdot)+\cos^{2}(\cdot)=1$). We also found it beneficial to have two separate networks, one predicting the position and the other predicting the rotation. In our implementation, for each of the network, we used a six-layer MLP with hidden dimension $128$.
@@
During testing, we need to predict $T$ steps into the future given a sequence of actions $a_{t:t+T}$. We do so by recursively applying the single-step predictor (A1). As we will show in §4, this simple MLP-based single-step predictor is sufficient for robust online planning. As introduced in §2, to ensure the dynamics of the world model is learned from a richer distribution than the expert demonstrations, the training dataset of the world model is expanded with random exploration demonstrations which attempt to record more pusher-object interactions at different contact points. To be specific, we let human operators (who are different from the operator of the expert demonstrations) collect $7$ long random play demonstrations and 1 special long demonstration that randomly collides the object with the boundary wall. In addition, we also generate $7$ sets of “noisy expert demonstrations” where the original expert actions are contaminated with different scales of Gaussian noise.
@@
Results. In Table 1, we study how different variations of $K$ and $M$ affect the performance. When $K=1$ and $M=0$, this is the baseline state-based diffusion policy trained from behavior cloning [7]. In GPC-RANK, if we increase the number of action proposals to $K=100$, the score increases to around 0.9, which is about $10\%$ increase from the baseline. Notably, due to parallelization inside GPC-RANK, the runtime does not increase when $K=100$. We also tried larger $K$, such as $1000$ and $5000$, and the performance further increases to $0.934$. Note that in the last column of Table 1, we provided a “performance upper bound” which is obtained by using the groundtruth simulator (instead of the learned world model) to rank action proposals at $K=5000$. For GPC-OPT, if we sample one action proposal and perform gradient ascent from it (e.g., doing $30$ gradient steps), the score can be increased to $0.886$, which is also quite significant compared to the baseline. Then in GPC-RANK+OPT, we tried different combinations of $K$ and $M$. The improvement remains substantial, indicating the robustness of GPC and shows both sampling multiple proposals and gradient-based optimization consistently improve the performance.
In the first two columns of Table 1, we present the results for another baseline of using Offline RL to rank and select actions. The Offline RL method follows that of [31] and is described in Appendix E. Crucially, the Offline RL method only uses the expert demonstrations and does not leverage the random exploration data (because the random exploration data do not have labeled rewards). We can see that the performance of Offline RL is similar or even worse than the case of behavior cloning. A plausible reason for this is that the Offline RL method does not explicitly seek to learn the dynamics from the random explorations and hence the learned value function may be misleading for ranking action proposals.
Finally, in Table 4, we present the results of using a world model trained without random exploration data. We observe that, when the world model is only trained from expert demonstrations, the performance of both GPC-RANK and GPC-OPT gets worse. This demonstrates the importance of using random exploration data to learn an accurate world model.
@@
Table 4: The importance of random exploration in world model learning (state-based Push-T).
Appendix D Vision-based Simulation Tasks
We test on four vision-based simulation tasks, named as Push-T, Triangle Drawing, Block Stacking, and Cube & Sphere Swapping.